# SimSort: A Data-Driven Framework for Spike Sorting by Large-Scale Electrophysiology Simulation

### v2

@@ A.1.2 Test datasets
Hybrid dataset. This dataset was composed of recordings provided by Spikeforest [41]. These recordings were generated by Kilosort, with waveform templates recorded at a 5 µm electrode spacing and  1/f background noise. It consisted of two subsets: 9 static tetrode recordings and 9 drift tetrode recordings that simulated sinusoidal probe movements. For each recording, we selected units with an SNR greater than 3 (unit IDs meeting this criterion were provided by SpikeForest to ensure a fair comparison), comprising 10–15 ground-truth units per recording. The dataset we prepared supported two modes of use. To evaluate overall spike sorting and detection performance, we used the raw data as input. For assessing identification performance, we constructed a waveform dataset by extracting 500 waveforms (60-timesteps length) per unit.
@@
where $\theta$ is the direction of the moving grating and $R(\theta)$ is the neuronal response to direction $\theta$. The imaginary unit is denoted as $i$ whose square is $-1$. The gOSI ranges from 0 to 1, with 0 indicating no orientation selectivity (equal response to all directions), and 1 indicating maximal selectivity (response only to a single orientation).
@@ A.2 Technical Details of Simulation
Intracellular Simulation. Intracellular electrophysiological activity was simulated using the NEURON package [30, 47]. Pink noise, scaled to the rheobase (the minimum current needed to elicit an action potential in a neuron), was injected into each neuron’s soma to evoke stochastic action potentials, simulating biologically realistic background noise. The temporal resolution was set to 0.1 ms, and simulations spanned a total duration of 600 seconds. Neurons were initialized with a resting membrane potential of -70 mV and maintained at a physiological temperature of 34°C. Synaptic mechanisms and biophysical properties were dynamically compiled and loaded to ensure compatibility with the NEURON simulation environment.
For each neuron, a multi-compartmental model was used to calculate transmembrane currents [48, 49, 50, 51]. The neuron was divided into compartments, each described as an equivalent electrical circuit. The dynamics of the membrane potential in each compartment $n$ were governed by Kirchhoff’s current law [52], where the sum of all currents entering or leaving a node must equal zero. Considering that extracellular potential changed much slower than ion channel dynamics, we assumed it constant. The temporal evolution of the membrane potential $V_{n}$ in compartment $n$ was given by $g_{n,n+1}(V_{n+1}-V_{n})-g_{n-1,n}(V_{n}-V_{n-1})=C_{n}\frac{dV_{n}}{dt}+\sum_{j}I_{n}^{(j)},$ where $g_{n,n+1}$ and $g_{n-1,n}$ represented the conductances between neighboring compartments, $C_{n}$ was the membrane capacitance, and $\sum_{j}I_{n}^{(j)}$ accounted for ionic currents and any externally applied currents.
@@
Extracellular Simulation. Extracellular potentials were modeled using the volume conductor theory, which links transmembrane currents to extracellular potentials recorded at electrode sites [53, 54, 55]. In each simulation trial, five neurons were randomly selected from the available models and positioned within a 100 × 100 × 100 $\mu$m3 cubic space. These neurons consisted of different types to represent the diversity of cortical microcircuits. A virtual tetrode electrode with four recording sites was placed randomly within the same space to capture extracellular potentials.
@@ A.3 Spike Detection
where $[t_{s},t_{e}]$ denoted the boundaries of a group of adjacent peaks.
@@
[…]
@@
where $[t_{s},t_{e}]$ defined the boundaries of a cluster of adjacent peaks. This step ensured that closely spaced peaks were consolidated into a single spike.
@@ A.5 Definitions of Metrics
To evaluate the performance of spike sorting algorithms, we adopted standard metrics as defined in the context of spike sorting and ground-truth comparisons [41, 44]. These metrics included accuracy, precision, and recall, which were derived based on the matching between the sorted spike train and the ground-truth spike train.
@@ Appendix B Supplementary Figures
Table S4: Performance comparison of spike identification on 50 test sets from IBL Neuropixels dataset. ARI values are presented as mean $\pm$ standard deviation.

### v3

@@ A.1.2 Test datasets
Hybrid dataset. This dataset was composed of recordings provided by Spikeforest [maglandSpikeForestReproducibleWebfacing2020]. These recordings were generated by Kilosort, with waveform templates recorded at a 5 µm electrode spacing and  1/f background noise. It consisted of two subsets: 9 static tetrode recordings and 9 drift tetrode recordings that simulated sinusoidal probe movements. For each recording, we selected units with an SNR greater than 3 (unit IDs meeting this criterion were provided by SpikeForest to ensure a fair comparison), comprising 10–15 ground-truth units per recording. The dataset we prepared supported two modes of use. To evaluate overall spike sorting and detection performance, we used the raw data as input. For assessing identification performance, we constructed a waveform dataset by extracting 500 waveforms (60-timesteps length) per unit.
@@
where $\theta$ is the direction of the moving grating and $R(\theta)$ is the neuronal response to direction $\theta$. The imaginary unit is denoted as $i$ whose square is 1-1. The gOSI ranges from 0 to 1, with 0 indicating no orientation selectivity (equal response to all directions), and 1 indicating maximal selectivity (response only to a single orientation).
@@ A.2 Technical Details of Simulation
Intracellular Simulation. Intracellular electrophysiological activity was simulated using the NEURON package [hinesNEURONSimulationEnvironment1997, hinesNeuronToolNeuroscientists2001]. Pink noise, scaled to the rheobase (the minimum current needed to elicit an action potential in a neuron), was injected into each neuron’s soma to evoke stochastic action potentials, simulating biologically realistic background noise. The temporal resolution was set to 0.1 ms, and simulations spanned a total duration of 600 seconds. Neurons were initialized with a resting membrane potential of -70 mV and maintained at a physiological temperature of 34°C. Synaptic mechanisms and biophysical properties were dynamically compiled and loaded to ensure compatibility with the NEURON simulation environment.
For each neuron, a multi-compartmental model [hayModelsNeocorticalLayer2011, markramReconstructionSimulationNeocortical2015, gouwensSystematicGenerationBiophysically2018, billehSystematicIntegrationStructural2020] was used to calculate transmembrane currents [lindenLFPyToolBiophysical2014, hagenViSAPyPythonTool2015, hagenMultimodalModelingNeural2018, buccinoMEArecFastCustomizable2021]. The neuron was divided into compartments, each described as an equivalent electrical circuit. The dynamics of the membrane potential in each compartment $n$ were governed by Kirchhoff’s current law [holmesPassiveCableModeling2009], where the sum of all currents entering or leaving a node must equal zero. Considering that extracellular potential changed much slower than ion channel dynamics, we assumed it constant. The temporal evolution of the membrane potential $V_{n}$ in compartment $n$ was given by $g_{n,n+1}(V_{n+1}-V_{n})-g_{n-1,n}(V_{n}-V_{n-1})=C_{n}\frac{dV_{n}}{dt}+\sum_{j}I_{n}^{(j)},$ where $g_{n,n+1}$ and $g_{n-1,n}$ represented the conductances between neighboring compartments, $C_{n}$ was the membrane capacitance, and $\sum_{j}I_{n}^{(j)}$ accounted for ionic currents and any externally applied currents.
@@
[…]
@@ A.3 Spike Detection
where $t_{s},t_{e}]$ denoted the boundaries of a group of adjacent peaks.
@@
where $Th_{\text{single\_ch}}$ was the amplitude threshold, and $r\in\{1,\dots,\text{loc\_range}\}$. These conditions ensured that the detected peaks were above the threshold and were local maxima within the range $r$. Peaks identified by this criterion were further validated within a broader range long_range to ensure they represented significant spike events.
@@
where $t_{s},t_{e}]$ defined the boundaries of a cluster of adjacent peaks. This step ensured that closely spaced peaks were consolidated into a single spike.
@@ A.5 Definitions of Metrics
To evaluate the performance of spike sorting algorithms, we adopted standard metrics as defined in the context of spike sorting and ground-truth comparisons [maglandSpikeForestReproducibleWebfacing2020, buccinoSpikeInterfaceUnifiedFramework2020]. These metrics included accuracy, precision, and recall, which were derived based on the matching between the sorted spike train and the ground-truth spike train.
@@ Appendix B Supplementary Figures
Table S4: Performance comparison of spike identification on 50 test sets from IBL Neuropixels dataset. ARI values are presented as mean $\pm$ standard deviation. Statistical significance was assessed with paired t-tests: $p=0.0012$ for SimSort vs. CEED, $p=0.0554$ for SimSort vs. CEED+UMAP.